Average speed of ball after second collision with ground dropped from 20 m

Grade:12th pass

1 Answers

Lakshay Kumar
6 years ago
Can you pls specify the value of coefficient of restitution ; "e". You can calculate it by following following relation v=e^n*v0=>v=e^n√2gh.Where n=after nth collisionHope you will understand it